OpenStudy (anonymous):

Rachel, Ahmad, and Dale sent a total of 107 text messages over their cell phones during the weekend. Dale sent 7 fewer messages than Rachel. Ahmad sent 2 times as many messages as Dale. How many messages did they each send?

OpenStudy (nikato):

Systems of equations D=dale A=Ahmad R=rachel

OpenStudy (anonymous):

dale 28. R=57 A=22 this is a rough guess.

OpenStudy (nikato):

So R+A+D=107 D=R-7 A=2D

OpenStudy (anonymous):

NVM